OpenStudy (helpmemath):

Find each commission. 6% on $3,000 in sales. Someone explain how to do this?

jimthompson5910 (jim_thompson5910):

6% of 3,000 = (6/100)*3000 6% of 3,000 = 0.06*3000 6% of 3,000 = ???

OpenStudy (helpmemath):

Thanks!

jimthompson5910 (jim_thompson5910):

np